One endpoint of a line segment is at (-2,4). The midpoint of the line segment is at (1,1). What is the coordinate of the second enpoint?
Answer by stanbon(50455)

One endpoint of a line segment is at (-2,4). The midpoint of the line segment is at (1,1). What is the coordinate of the second enpoint?
---
Each coordinate of the midpoint is the average of
the corresponding coordinates of the end points.
---
Let the missing endpoint be (x,y)
Equations:
(x + (-2))/2 = 1
x-2 = 2
x = 4
----
(y+4)/2 = 1
y+4 = 2
y = -2
-----
2nd endpoint (4,-2)
===========================
Cheers,
Stan H.